# Detect Anything 3D in the Wild

### v2

@@ Abstract
Despite the success of deep learning in close-set 3D object detection, existing approaches struggle with zero-shot generalization to novel objects and camera configurations. We introduce DetAny3D, a promptable 3D detection foundation model capable of detecting any novel object under arbitrary camera configurations using only monocular inputs. Training a foundation model for 3D detection is fundamentally constrained by the limited availability of annotated 3D data, which motivates DetAny3D to leverage the rich prior knowledge embedded in extensively pre-trained 2D foundation models to compensate for this scarcity. To effectively transfer 2D knowledge to 3D, DetAny3D incorporates two core modules: the 2D Aggregator, which aligns features from different 2D foundation models, and the 3D Interpreter with Zero-Embedding Mapping, which mitigates catastrophic forgetting in 2D-to-3D knowledge transfer. Experimental results validate the strong generalization of our DetAny3D, which not only achieves state-of-the-art performance on unseen categories and novel camera configurations, but also surpasses most competitors on in-domain data. DetAny3D sheds light on the potential of the 3D foundation model for diverse applications in real-world scenarios, e.g., rare object detection in autonomous driving, and demonstrates promise for further exploration of 3D-centric tasks in open-world settings. More visualization results can be found at our code repository.
@@ 1 Introduction
3D object detection is a fundamental technology for autonomous systems [36, 15, 14, 49, 48, 12], robotics [84, 6, 67], and augmented reality [43, 52]. 3D perception not only enables machines to perceive and interact with the physical world, but also serves as a foundational input for more advanced tasks, such as behavior decision [20, 3, 11, 31], world modeling [22, 23, 38] and 3D scene reconstruction [50, 73, 75]. For practical deployment, a generalizable 3D detector ideally should detect arbitrary objects from easily accessible inputs, such as monocular images, without relying on specific sensor parameters. Such a model would be highly adaptable and reliable for various downstream tasks in diverse and unpredictable environments [36, 15, 43, 84]. Also, accurate detection results provided by such a detector (e.g., generating 3D bounding boxes for even images from the internet) make it a versatile tool, paving the way for scalable 3D systems that leverage Internet-scale data and advance toward open-world scenarios [22, 23, 50, 73, 38].
@@
Previous research, exemplified by Omni3D [8], has attempted to improve the generalization of the 3D detection system through multi-dataset training [68, 35, 8, 40]. However, despite utilizing large datasets to train a unified detector [8, 40], these approaches provide limited generalization to novel camera configurations and cannot detect unseen object categories beyond predefined label spaces. Therefore, developing a 3D detection foundation model with strong zero-shot generalizability, which is capable of detecting any unseen object under arbitrary camera configurations, remains a crucial and unsolved problem.
@@
While recent advances in 2D foundation models [33, 51, 56, 44] demonstrate remarkable zero-shot capabilities. Segment Anything Model (SAM) [33] features a promptable inference mechanism, supporting user-friendly prompts like points and boxes to segment user-specified objects. Their impressive generalization ability stems from training on billions of annotated images. However, in 3D object detection, the available labeled data is limited to only millions of samples—typically 3-4 orders of magnitude smaller than in 2D images. Such severe data scarcity [86, 74] poses a fundamental challenge, making it nearly infeasible to train a 3D foundation model from scratch.
@@
In this work, we present DetAny3D, a promptable 3D detection foundation model designed for generalizable 3D object detection using only monocular images (see Figure 1). Given the inherent scarcity of 3D annotated data, we achieve strong generalization from two critical perspectives: model architecture and data utilization. The central insight of our approach is to leverage the extensive prior knowledge encoded within two broadly pre-trained 2D foundation models—SAM [33] and DINO [51, 10]—thus unlocking effective zero-shot 3D detection capabilities with minimal available 3D data.
@@
Specifically, we adopt SAM as our promptable backbone, capitalizing on its versatile and robust object understanding capability derived from large-scale 2D data. Concurrently, we utilize DINO [51] depth-pretrained by UniDepth [54], to offer redundant 3D geometric priors [76, 7], which plays a pivotal role for accurate 3D detection in a monocular setting. To integrate the complementary features from SAM and DINO more effectively, we propose the 2D Aggregator, an attention-based mechanism that aligns these features and dynamically optimizes their contributions via learnable gating. 2D Aggregator fully exploits the strengths of each foundation model.
@@
To further address the challenge of effectively transferring knowledge from 2D to 3D, we introduce the 3D Interpreter. Central to the 3D Interpreter is the Zero-Embedding Mapping (ZEM) mechanism, which mitigates catastrophic forgetting issues common in cross-domain learning. By stabilizing the training process across diverse datasets with varying camera parameters, scene complexities, and depth distributions, the ZEM mechanism enables progressive zero-shot 3D grounding capabilities, significantly enhancing model generalization.
@@
  • DetAny3D introduces 2D Aggregator to effectively fuse the features from two 2D foundation models SAM and depth-pretrained DINO, which provides pivot shape and 3D geometric priors for vaious objects, respectively.
  • In 2D-to-3D knowledge transfer, DetAny3D involves Zero-Embedding Mapping in 3D Interpreter to address the catastrophic forgetting dilemma, enabling the model to train stably across datasets with diverse camera parameters, varying scenes, and different depth distributions.
@@ 2 Related works
Existing 3D object detection systems have predominantly focused on single-dataset optimization, achieving strong performance on benchmark datasets like KITTI [24] and nuScenes [9] through task-specific architectures [39, 80, 66, 14, 45, 18, 42, 41]. While effective in constrained scenarios, these approaches suffer from significant domain gaps when deployed in new contexts, primarily due to their reliance on limited sensor-specific data and closed-set assumptions. Recent works, exemplified by Omni3D [8], have demonstrated the potential of multi-dataset training. Models like Cube R-CNN [8] and UniMODE [40] train a universal monocular 3D detector across multiple datasets, achieving some level of robustness to camera parameters, but are still restricted to predefined classes. V-MIND [32] further addresses the data scarcity challenge by generating pseudo 3D training data from large-scale 2D annotations. Towards more general detection, OV-Uni3DETR [69] pioneers open-set detection that is able to detect with multimodal inputs, but it is trained separately for indoor and outdoor domains, thereby limiting its overall generalization. More recently, OVMono3D [74] leverages grounding DINO’s [44] 2D results with a 3D head on unified datasets. However, it does not fully exploit the priors contained in 2D foundation models, leading to performance constraints tied to the limited 3D data. In contrast, our approach fully capitalizes on the knowledge distilled in 2D foundation models while leveraging abundant 3D-related data, thereby enabling the detection of any 3D object from arbitrary monocular inputs.
@@ 2.2 Vision Foundation Models
Foundation models have demonstrated significant potential across various domains. For example, language foundation models such as GPT-4 [1] and DeepSeek [5, 26], trained on massive internet-scale corpora, have achieved impressive capabilities in natural language processing across diverse fields [63, 1, 5, 60, 81, 82]. Similarly, foundation models in the vision domain have made remarkable strides [51, 33, 44, 56, 37, 79, 29]. DINOv2 [51], trained on a vast range of curated data from diverse sources, is capable of producing general-purpose visual features that work seamlessly across different image distributions and tasks. SAM [33] has taken a step further in the vision domain by introducing promptability, enabling models to generalize to novel visual concepts through large-scale data training and continuous model refinement. In recent years, the development of foundation models in the 3D domain has started to take initial steps [55, 13, 85, 83, 78, 28]. Most existing 3D foundation models are often combined with vision-language models (VLMs) [55, 13, 85, 27], relying on point clouds as input to help the language models understand 3D [13, 85]. While these methods are valuable for scene understanding and semantic tasks, they do not directly provide precise 3D detection results. Moreover, point cloud inputs significantly restrict the use cases [72], as they are not always accessible in many practical scenarios. In contrast to these approaches, we aim to develop a foundation model specifically dedicated to 3D detection tasks with the most general inputs, monocular images. By leveraging the powerful priors from 2D vision foundation models, our approach enables the detection of any 3D object with arbitrary camera configurations, presenting a broad range of practical applications.
@@ 3.1 Overview
Specifically, the monocular image is embedded in parallel by two foundational models: SAM [33] for low-level pixel information, underpins the entire promptable architecture. And depth-pretrained DINO [54, 51], which provide rich high-level geometric knowledge, excels in depth-related tasks. These complementary 2D features are then fused through our proposed 2D Aggregator (see Figure 2(b)), which hierarchically aligns low-level and high-level information using cross-attention layers. The fused features are subsequently passed to the Depth/Camera Module, which extracts the camera and camera-aware depth embedding, collectively referred to as geometric embedding.
The geometric embedding and the 3D bounding box token with encoded prompt tokens are then fed into the 3D Interpreter (see Figure 2(c)), which employs a structure similar to the SAM decoder along with a specialized Zero-Embedding Mapping (ZEM) mechanism. 3D Interpreter injects 3D geometric features while preventing the catastrophic forgetting dilemma in 2D-to-3D knowledge transfer, achieving progressive 3D grounding. Finally, the model predicts 3D boxes based on the hidden states of the 3D box token. Our DetAny3D is trained on selected seen classes and can detect any unseen classes in a zero-shot manner.
@@
Feature Extraction. Given an input image, the SAM encoder extracts high-resolution spatial features $\mathbf{F}_{s}\in\mathbb{R}^{H_{s}\times W_{s}\times C},$ capturing fine-grained details and boundaries. Simultaneously, the DINO encoder outputs geometry-aware embeddings $\mathbf{F}_{d}\in\mathbb{R}^{H_{d}\times W_{d}\times C},$ which is depth-pretrained by Unidepth [54] and provide robust priors for depth and intrinsics. Following the design of ViT Adapter [16], we also employ a convolutional structure to produce preliminary image features, denoted as $\mathbf{F}_{q}^{0},$ serving as the initial query for subsequent attention-based fusion.
@@
Geometric Embeddings. The fused features $\hat{\mathbf{F}}_{\text{fused}}^{i},$ $i\in[1,2,3,4],$ are subsequently processed by the depth and camera modules, following the Unidepth [54] architecture. Specifically, these modules predict the camera embedding $\mathbf{C}$ and camera-aware depth embedding $\mathbf{D|C}$, referred as the geometric embedding $\mathbf{G}=\{\mathbf{D|C},\,\mathbf{C}\}$. These modules provide aligned depth and camera parameters under the monocular depth ill-posed problem. Further details can be found in Section 7.1.
@@
The diverse 3D object supervisions across various scenarios, depths, and camera intrinsics introduce challenges to model training. Our 3D Interpreter aims to progressively integrate geometric information while preventing catastrophic forgetting in 2D-to-3D knowledge transfer. We introduce Zero-Embedding Mapping (ZEM) mechanism, which incrementally infuses 3D geometry into the decoder via zero-initialized layers—without disrupting the original 2D features. As Figure 2(c) shows, the 3D Interpreter comprises three main components: the Two-Way Transformer, the Geometric Transformer, and the 3D bounding box heads.
Two-Way Transformer. Following the SAM design, we first concatenate the 3D bounding box token with prompt-related tokens to form the query:
@@
where $\mathbf{T}_{\mathrm{3D},i}$ denotes the 3D bounding box token for the $i$-th object, $\mathbf{T}_{\mathrm{p},i}$ is the prompt-related token, and [;][$\cdot$;\,$\cdot$][ ⋅ ; ⋅ ] denotes vector concatenation. The SAM encoder output $\mathbf{F}_{s}$ serves as both key and value for the first Two-Way Transformer layer, yielding:
@@
3D Bounding Box Heads. Finally, $\mathbf{O}$ is fed into the 3D bounding box heads to calculate the final predictions, which follows typical architectures from standard 3D detection frameworks [80, 8, 66]: $B_{\textrm{3D}}(x,\,y,\,z,\,w,\,h,\,l,\,R,\,S\,)$ where $x,y,z$ specify the 3D box center, $w,h,l$ are its dimensions, $R$ is the rotation matrix, and $S$ is the predicted 3D Intersection over Union (IoU) score.

### v3

@@ Abstract
Despite the success of deep learning in close-set 3D object detection, existing approaches struggle with zero-shot generalization to novel objects and camera configurations. We introduce DetAny3D, a promptable 3D detection foundation model capable of detecting any novel object under arbitrary camera configurations using only monocular inputs. Training a foundation model for 3D detection is fundamentally constrained by the limited availability of annotated 3D data, which motivates DetAny3D to leverage the rich prior knowledge embedded in extensively pre-trained 2D foundation models to compensate for this scarcity. To effectively transfer 2D knowledge to 3D, DetAny3D incorporates two core modules: the 2D Aggregator, which aligns features from different 2D foundation models, and the 3D Interpreter with Zero-Embedding Mapping, which stabilizes early training in 2D-to-3D knowledge transfer. Experimental results validate the strong generalization of our DetAny3D, which not only achieves state-of-the-art performance on unseen categories and novel camera configurations, but also surpasses most competitors on in-domain data. DetAny3D sheds light on the potential of the 3D foundation model for diverse applications in real-world scenarios, e.g., rare object detection in autonomous driving, and demonstrates promise for further exploration of 3D-centric tasks in open-world settings. More visualization results can be found at our code repository.
@@ 1 Introduction
3D object detection is a fundamental technology for autonomous systems [li2019gs3d, chen2017multi, chen2016monocular, mao20233d, ma20233d, chen2024end], robotics [zhu2014single, biegelbauer2007efficient, wang2024embodiedscan], and augmented reality [liu2019edge, park2008multiple]. 3D perception not only enables machines to perceive and interact with the physical world, but also serves as a foundational input for more advanced tasks, such as behavior decision [fang2020graspnet, asif2018graspnet, casas2021mp3, hu2023planning], world modeling [gao2023magicdrive, gao2024magicdrive3d, li2024drivingdiffusion] and 3D scene reconstruction [nie2020total3dunderstanding, yang2025scenecraft, yao2025cast]. For practical deployment, a generalizable 3D detector ideally should detect arbitrary objects from easily accessible inputs, such as monocular images, without relying on specific sensor parameters. Such a model would be highly adaptable and reliable for various downstream tasks in diverse and unpredictable environments [li2019gs3d, chen2017multi, liu2019edge, zhu2014single]. Also, accurate detection results provided by such a detector (e.g., generating 3D bounding boxes for even images from the internet) make it a versatile tool, paving the way for scalable 3D systems that leverage Internet-scale data and advance toward open-world scenarios [gao2023magicdrive, gao2024magicdrive3d, nie2020total3dunderstanding, yang2025scenecraft, li2024drivingdiffusion].
@@
Previous research, exemplified by Omni3D [brazil2023omni3d], has attempted to improve the generalization of the 3D detection system through multi-dataset training [wang2023uni3detr, kolodiazhnyi2024unidet3d, brazil2023omni3d, li2024unimode]. However, despite utilizing large datasets to train a unified detector [brazil2023omni3d, li2024unimode], these approaches provide limited generalization to novel camera configurations and cannot detect unseen object categories beyond predefined label spaces. Therefore, developing a 3D detection foundation model with strong zero-shot generalizability, which is capable of detecting any unseen object under arbitrary camera configurations, remains a crucial and unsolved problem.
@@
While recent advances in 2D foundation models [kirillov2023segment, oquab2024dinov2, radford2021learning, liu2024grounding] demonstrate remarkable zero-shot capabilities. Segment Anything Model (SAM) [kirillov2023segment] features a promptable inference mechanism, supporting user-friendly prompts like points and boxes to segment user-specified objects. Their impressive generalization ability stems from training on billions of annotated images. However, in 3D object detection, the available labeled data is limited to only millions of samples—typically 3-4 orders of magnitude smaller than in 2D images. Such severe data scarcity [zuo2024towards, yao2024open] poses a fundamental challenge, making it nearly infeasible to train a 3D foundation model from scratch.
@@
In this work, we present DetAny3D, a promptable 3D detection foundation model designed for generalizable 3D object detection using only monocular images (see Figure 1). Given the inherent scarcity of 3D annotated data, we achieve strong generalization from two critical perspectives: model architecture and data utilization. The central insight of our approach is to leverage the extensive prior knowledge encoded within two broadly pre-trained 2D foundation models—SAM [kirillov2023segment] and DINO [oquab2024dinov2, caron2021emerging]—thus unlocking effective zero-shot 3D detection capabilities with minimal available 3D data.
@@
Specifically, we adopt SAM as our promptable backbone, capitalizing on its versatile and robust object understanding capability derived from large-scale 2D data. Concurrently, we utilize DINO [oquab2024dinov2] depth-pretrained by UniDepth [piccinelli2024unidepth], to offer redundant 3D geometric priors [yin2023metric3d, bochkovskii2024depth], which plays a pivotal role for accurate 3D detection in a monocular setting. To integrate the complementary features from SAM and DINO more effectively, we propose the 2D Aggregator, an attention-based mechanism that aligns these features and dynamically optimizes their contributions via learnable gating. 2D Aggregator fully exploits the strengths of each foundation model.
@@
To further address the challenge of effectively transferring knowledge from 2D to 3D, we introduce the 3D Interpreter. Central to the 3D Interpreter is the Zero-Embedding Mapping (ZEM) mechanism, which ensures stable 2D-to-3D mapping by reducing early-stage interference and preserving pretrained 2D priors. By stabilizing the training process across diverse datasets with varying camera parameters, scene complexities, and depth distributions, the ZEM mechanism enables progressive zero-shot 3D grounding capabilities, significantly enhancing model generalization.
@@
  • DetAny3D introduces 2D Aggregator to effectively fuse the features from two 2D foundation models SAM and depth-pretrained DINO, which provide pivot shape and 3D geometric priors for various objects, respectively.
  • In 2D-to-3D knowledge transfer, DetAny3D involves Zero-Embedding Mapping in 3D Interpreter to achieve stable 2D-to-3D mapping, enabling the model to train robustly across datasets with diverse camera parameters, varying scenes, and different depth distributions.
@@ 2 Related works
Existing 3D object detection systems have predominantly focused on single-dataset optimization, achieving strong performance on benchmark datasets like KITTI [geiger2013vision] and nuScenes [caesar2020nuscenes] through task-specific architectures [li2024bevformer, zhang2023monodetr, wang2021fcos3d, chen2016monocular, liu2020smoke, dasgupta2016delay, lin2022sparse4d, liang2022bevfusion]. While effective in constrained scenarios, these approaches suffer from significant domain gaps when deployed in new contexts, primarily due to their reliance on limited sensor-specific data and closed-set assumptions. Recent works, exemplified by Omni3D [brazil2023omni3d], have demonstrated the potential of multi-dataset training. Models like Cube R-CNN [brazil2023omni3d] and UniMODE [li2024unimode] train a universal monocular 3D detector across multiple datasets, achieving some level of robustness to camera parameters, but are still restricted to predefined classes. V-MIND [jhang2025vmind] further addresses the data scarcity challenge by generating pseudo 3D training data from large-scale 2D annotations. Towards more general detection, OV-Uni3DETR [wang2024ov] pioneers open-set detection that is able to detect with multimodal inputs, but it is trained separately for indoor and outdoor domains, thereby limiting its overall generalization. More recently, OVMono3D [yao2024open] leverages Grounding DINO’s [liu2024grounding] 2D results with a 3D head on unified datasets. However, it does not fully exploit the priors contained in 2D foundation models, leading to performance constraints tied to the limited 3D data. In contrast, our approach fully capitalizes on the knowledge distilled in 2D foundation models while leveraging abundant 3D-related data, thereby enabling the detection of any 3D object from arbitrary monocular inputs.
@@ 2.2 Vision Foundation Models
Foundation models have demonstrated significant potential across various domains. For example, language foundation models such as GPT-4 [achiam2023gpt] and DeepSeek [bi2024deepseek, guo2024deepseek], trained on massive internet-scale corpora, have achieved impressive capabilities in natural language processing across diverse fields [touvron2023llama, achiam2023gpt, bi2024deepseek, sima2024drivelm, zhang2024llama, zhang2024mavis]. Similarly, foundation models in the vision domain have made remarkable strides [oquab2024dinov2, kirillov2023segment, liu2024grounding, radford2021learning, li2022grounded, zhang2023personalize, guo2025can]. DINOv2 [oquab2024dinov2], trained on a vast range of curated data from diverse sources, is capable of producing general-purpose visual features that work seamlessly across different image distributions and tasks. SAM [kirillov2023segment] has taken a step further in the vision domain by introducing promptability, enabling models to generalize to novel visual concepts through large-scale data training and continuous model refinement. In recent years, the development of foundation models in the 3D domain has started to take initial steps [qi2025gpt4scene, chen2024ll3da, zhu2024unifying, zhu2023ponderv2, zhang2022pointclip, guo2024sam2point]. Most existing 3D foundation models are often combined with vision-language models (VLMs) [qi2025gpt4scene, chen2024ll3da, zhu2024unifying, guo2023point], relying on point clouds as input to help the language models understand 3D [chen2024ll3da, zhu2024unifying]. While these methods are valuable for scene understanding and semantic tasks, they do not directly provide precise 3D detection results. Moreover, point cloud inputs significantly restrict the use cases [Yang_2024_CVPR], as they are not always accessible in many practical scenarios. In contrast to these approaches, we aim to develop a foundation model specifically dedicated to 3D detection tasks with the most general inputs, monocular images. By leveraging the powerful priors from 2D vision foundation models, our approach enables the detection of any 3D object with arbitrary camera configurations, presenting a broad range of practical applications.
@@ 3.1 Overview
Specifically, the monocular image is embedded in parallel by two foundational models: SAM [kirillov2023segment] for low-level pixel information, underpins the entire promptable architecture. And depth-pretrained DINO [piccinelli2024unidepth, oquab2024dinov2], which provide rich high-level geometric knowledge, excels in depth-related tasks. These complementary 2D features are then fused through our proposed 2D Aggregator (see Figure 2(b)), which hierarchically aligns low-level and high-level information using cross-attention layers. The fused features are subsequently passed to the Depth/Camera Module, which extracts the camera and camera-aware depth embedding, collectively referred to as geometric embedding.
The geometric embedding and the 3D bounding box tokens with encoded prompt tokens are then fed into the 3D Interpreter (see Figure 2(c)), which employs a structure similar to the SAM decoder along with a specialized Zero-Embedding Mapping (ZEM) mechanism. 3D Interpreter injects 3D geometric features while ensuring stable 2D-to-3D knowledge transfer, enabling progressive 3D grounding across diverse data domains. Finally, the model predicts 3D boxes based on the hidden states of the 3D box tokens. Our DetAny3D is trained on selected seen classes and can detect any unseen classes in a zero-shot manner.
@@
Feature Extraction. Given an input image, the SAM encoder extracts high-resolution spatial features $\mathbf{F}_{s}\in\mathbb{R}^{H_{s}\times W_{s}\times C},$ capturing fine-grained details and boundaries. Simultaneously, the DINO encoder outputs geometry-aware embeddings $\mathbf{F}_{d}\in\mathbb{R}^{H_{d}\times W_{d}\times C},$ which is depth-pretrained by UniDepth [piccinelli2024unidepth] and provides robust priors for depth and intrinsics. Following the design of ViT Adapter [chenvision], we also employ a convolutional structure to produce preliminary image features, denoted as $\mathbf{F}_{q}^{0},$ serving as the initial query for subsequent attention-based fusion.
@@
Geometric Embeddings. The fused features $\hat{\mathbf{F}}_{\text{fused}}^{i},$ $i\in[1,2,3,4],$ are subsequently processed by the depth and camera modules, following the UniDepth [piccinelli2024unidepth] architecture. Specifically, these modules predict the camera embedding $\mathbf{C}$ and camera-aware depth embedding $\mathbf{D|C}$, referred as the geometric embedding $\mathbf{G}=\{\mathbf{D|C},\,\mathbf{C}\}$. These modules provide aligned depth and camera parameters under the monocular depth ill-posed problem. Further details can be found in the Supplementary material Section 7.1.
@@
The diverse 3D object supervisions across various scenarios, depths, and camera intrinsics introduce challenges to model training. Our 3D Interpreter aims to progressively integrate geometric information while ensuring stable 2D-to-3D knowledge transfer. We introduce Zero-Embedding Mapping (ZEM) mechanism, which incrementally infuses 3D geometry into the decoder via zero-initialized layers—without disrupting the original 2D features. As Figure 2(c) shows, the 3D Interpreter comprises three main components: the Two-Way Transformer, the Geometric Transformer, and the 3D bounding box heads.
Two-Way Transformer. Following the SAM design, we first concatenate the 3D bounding box tokens with prompt-related tokens to form the query:
@@
where $\mathbf{T}_{\mathrm{3D},i}$ denotes the 3D bounding box token for the $i$-th object, $\mathbf{T}_{\mathrm{p},i}$ is the prompt-related token, and $\cdot$$\cdot$] denotes vector concatenation. The SAM encoder output $\mathbf{F}_{s}$ serves as both key and value for the first Two-Way Transformer layer, yielding:
@@
3D Bounding Box Heads. Finally, $\mathbf{O}$ is fed into the 3D bounding box heads to calculate the final predictions, which follows typical architectures from standard 3D detection frameworks [zhang2023monodetr, brazil2023omni3d, wang2021fcos3d]: $B_{\textrm{3D}}(x,\,y,\,z,\,w,\,h,\,l,\,R,\,S\,)$ where $x,y,z$ specify the 3D box center, $w,h,l$ are its dimensions, $R$ is the rotation matrix, and $S$ is the predicted 3D Intersection over Union (IoU) score.